singular and general solution 

e^3x(p-1)+p^3e^2y=0;X=e^x ,Y=e^y by substitution


1
Expert's answer
2021-06-28T16:29:02-0400
"e^{3x}(p-1)+p^3e^{2y}=0"

"e^y(1-p)=p^3e^{3(y-x)}"

Use the substitution "X=e^x, Y=e^y"


"dX=e^xdx, dY=e^y dy"

"\dfrac{dy}{dx}=\dfrac{e^x}{e^y}\cdot\dfrac{dY}{dX}=>p=\dfrac{X}{Y}\cdot\dfrac{dY}{dX}"

"Y(1-\dfrac{X}{Y}\cdot\dfrac{dY}{dX})=(\dfrac{X}{Y}\cdot\dfrac{dY}{dX}\cdot\dfrac{Y}{X})^3"

"Y-X\dfrac{dY}{dX}=(\dfrac{dY}{dX})^3"

The Clairaut equation 

"Y=XP+P^3, P=\dfrac{dY}{dX}"

Differentiating in "X," we have


"Y=XP+P^3"

"dY=XdP+PdX+3P^2dP"

Replace "dY" with "PdX" to obtain:


"PdX=XdP+PdX+3P^2dP"

"dP(X+3P^2)=0"

By equating the first factor to zero, we have


"dP=0=>P=C"

Now we substitute this into the differential equation.

The general solution is given by


"Y=CX+C^3, C\text{ is arbitrary constant}"

By equating the second term to zero we find that


"X+3P^2=0=>X=-3P^2"

This gives us the singular solution of the differential equation in parametric form:


"X=-3P^2"

"Y=XP+P^3"

By eliminating "P" from this system, we get the equation of the integral curve:


"P=\pm\sqrt{\dfrac{-X}{3}}, Y=\pm\dfrac{2}{3}X\sqrt{\dfrac{-X}{3}}"

Since "X=e^x, Y=e^y," then the singular solution of the differential equation does not exist.


The general solution is given by


"e^y=Ce^x+C^3, C\text{ is arbitrary constant}"